Nitrogen and hydrogen gases react to form ammonia gas as follows:
N₂(g) + 3 H₂(g) = 2 NH₃(g)
At a certain temperature and pressure, 1.2 L of N₂ reacts with 3.6 L of H₂. If all the N₂ and H₂ are consumed, what volume of NH₃, at the same temperature and pressure, will be produced?
1
Expert's answer
2020-11-17T10:33:19-0500

calculate the amount of substance N2 and H2

n(N2) = V(N2)/VM = 1,2 L / 22,4 L/moll = 0,054 moll

n(H2) = V(H2)/VM = 3,6 L / 22,4 L/moll = 0.16 moll


n(N2):n(H2) = 0.054:0.16 = 1:3


volume of NH3 can be calculated by both volume N2 or H2


If by both volume N2

V(NH3) = V(N2) × 2VM / VM = 1.2 L × (2×22.4 L/moll) / 22.4 L/moll = 2.4 L


If by both volume H2

V(NH3) = V(H2) × 2VM / 3VM = 3.6 L × (2×22.4 L/moll) / (3×22.4 L/moll) = 2.4 L



Volume of NH3 is 2.4 L